OpenStudy (anonymous):

Please check my work in the comments During the period 1985-2012, the projected enrollment B (in thousands of students) in public schools and the projected enrollment R in (in thousands of students) in private schools can be modeled by B = -18.53t^2 + 975.8t + 48,140 and R = 80.8t + 8,049 where t is the number of years. A: Write an equation that models the difference in the projected enrollments for public schools and private schools as a function of the number of years since 1985 B: Find the difference in projected enrollments for public schools and private schools in 2005.

OpenStudy (kropot72):

A: You need to find B - R. Subtract as follows: \[\large B-R=-18.53t^{2}+975.8t+48,140-(80.8t+8,049)\]

OpenStudy (kropot72):

@bookish610 Can you do the subtraction?

OpenStudy (anonymous):

yep doing it now. (on paper)

OpenStudy (anonymous):

would it be this? \[B-R = -99.33t + 955.8t + 40.091\]

OpenStudy (kropot72):

Not really. Perhaps the subtraction will be clearer if like terms are grouped, as follows: \[\large B-R=-18.53t^{2}+(975.8t-80.8t)+(48,140-8,049)=?\]

OpenStudy (kropot72):

Note that the term -18.53t^2 will be unchanged in the result of subtraction.

OpenStudy (anonymous):

\[B-R = -18.53t^{2} + 895t + 40,091\] ?

OpenStudy (kropot72):

Yes, that is correct. B: Now you need to find the number of years between 2005 and 1985, and then substitute that number for t in your equation for A:

OpenStudy (kropot72):

\[\large B-R(at\ 2005)=(-18.53\times20^{2})+(895\times20)+40,091=?\]

OpenStudy (anonymous):

\[B-R = 50,579\] ?

OpenStudy (anonymous):

@kropot72 Did I get it right?

OpenStudy (anonymous):

Can someone please tell me if I got this right?

OpenStudy (anonymous):

please guys

OpenStudy (-*.*-):



OpenStudy (anonymous):

I AM A STRESSED OUT HIGH SCHOOL STUDENT AND I DO NOT NEED YOUR SHENANIGANS

OpenStudy (anonymous):

I need someones help

jaynator495 (jaynator495):

x'D im sorry i just cant help but find that funny XD

OpenStudy (-*.*-):

medal for medal?anyone

OpenStudy (anonymous):

well i'm glad someone is finding humor in my PAIN

OpenStudy (-*.*-):

HAHAHAHAH

OpenStudy (kayders1997):

first t=20 so first plug in 20 for all the t's because 2005-195 is 20 years

OpenStudy (kayders1997):

1985

OpenStudy (kayders1997):

Does that make sense?

OpenStudy (anonymous):

Yeah. I think I did some of this earlier. Did you see the equation and what I think is the correct answer?

OpenStudy (kayders1997):

Just a second gotta grab my calculator :)

OpenStudy (anonymous):

ok

OpenStudy (kayders1997):

That is not what I got

OpenStudy (kayders1997):

First let's separate everything up so you have the equation -18.53t^2+975.8t+48140 so first let's plug in our 20 for the t's, and I like to put the parenthesis around the t's or the number your plugging in

OpenStudy (kayders1997):

So -18.53(20)^2+975.8(20)+48140 what do you get there?

OpenStudy (anonymous):

wait is that last part right? isn't it -18.53(20)^2+975.8(20)+40091 ?

OpenStudy (kayders1997):

It's 48140

OpenStudy (kayders1997):

At least that was what was typed

OpenStudy (kayders1997):

It that what it says into your book or wherever this problem is coming from?

OpenStudy (anonymous):

it's up farther where somebody was helping me figure out what the equation for part A was

OpenStudy (kayders1997):

Oh I see you guys reduced it okay you don't have to do the long process I was going to do but either way works too :) or should work

OpenStudy (anonymous):

so if I use the equation from earlier it comes out as 52,195

OpenStudy (kayders1997):

Yes that is right, I must have punched something In wrong my bad

OpenStudy (kayders1997):

You could also you the original equation and go (-18.53(20)^2975.8(20)+48140)-(80.8(20)+8049) doesn't matter which way you do it :) you get the same answer

OpenStudy (kayders1997):

Oops I messed up with the first part (-18.53(20)^2+975.8(20)+48140) but good job :)

OpenStudy (kayders1997):

Do you have any other questions

OpenStudy (anonymous):

yeah. three more two part word problems and one semi regular problem

OpenStudy (kayders1997):

Okay :) and sorry I didn't help you on Friday I got really busy

OpenStudy (kropot72):

Yes, your answer 50,579 is correct.